OpenStudy (anonymous):

find the lengths of both circular arts of the unit circle connecting the point (1,0) and the endpoint of the radius that makes an angle of 2.95 radians with the positive horizontal axis.

OpenStudy (hoblos):

2.95 radians = 2.95*180/(pi) degrees = 169 degrees so the first angle between the two points is 169° the lengths of first circular arc is 169° the lengths of second circular arc is 360-169=191°

OpenStudy (anonymous):

didn;t work though.

OpenStudy (anonymous):

wait a sec, what are 169 and 191 in terms of radians.

OpenStudy (hoblos):

169° = 2.95 rad 191° = 3.33 rad

OpenStudy (anonymous):

yep, that was what i though.... thanks man.

OpenStudy (hoblos):

yw